Drew Corp. Designs and manufactures mascot uniforms for high school, college, and professional sports teams. Since each team's uniform is unique in color and design, Drew uses a job order costing system. On January 1, the T-accounts for some of Drew's primary balance sheet accounts were as follows:

Raw materials inventory- 59400

Finished Goods Inventory-34600

Accounts receivable-47300

Work in process Inventory-25500

Cash-38200

accounts payable- credit 41900

During the year, the following events occurred:

1. Drew purchased raw materials costing 114000 on account.

2. Drew used 149300 of raw materials in production. Of these, 80% were classified as direct materials and 20% as indirect materials. (Drew maintains a single Raw Materials Inventory Account.)

3. Drew used 31200 hours of direct labor. The company's average direct labor rate was 11 per hour (credit Wages Payable).

4. The company's indirect labor cost was 168000 (credit Wages Payable).

5. Other manufacturing overhead costs the company incurred on account totaled 98400

6. Drew applied 316800 in manufacturing overhead

7. the company completed production of goods costing 795400

8. the company's cost of goods sold balance was 801300, before adjusting for over- or underapplied overhead.

9. Sales revenue was 1050000 (all sales were made on account).

10. Drew collected 803200

11. The company paid accounts payable of 201300

12. At year-end, all wages earned during the year had been paid.

A.) Calculate under- or overapplied overhead for the year.

B.)Assuming that Unresolved closes under- or overapplied overhead to Cost of Goods Sold, calculate the cost of goods sold for the year.

C.)Assuming that Unresolved prorates under- or overapplied overhead to the appropriate accounts, calculate the adjusted Work in Process Inventory, Finished Goods Inventory, and Cost of Goods Sold balances for the year
Answer #1
A) Actual manufacturing overhead:
$
Indirect materials (149300*20%) 29860
Indirect labor cost 168000
Other manufacturing overhead costs 98400
Total 296260
Manufacturing overhead applied=$ 316800
Manufacturing overhead applied > Actual manufacturing overhead
Hence,Manufacturing overhead is over-applied
Over-applied overehad=316800-296260=$ 20540
B) Over-applied overehead reduces the cost of goods sold by $ 20540.
Cost of goods sold=801300-20540=$ 780760
C) Over-applied overhead is prorated to the work in Process Inventory,finished Goods Inventory, and cost of Goods Sold based on the ending balance (unadjusted) of the appropriate accounts
Ending balances (unadjusted):
Work in process inventory
$
Beginning balance 25500
Add:
Direct materials cost (149300*80%) 119440
Direct labor cost (31200*11) 343200
Manufacturing overhead applied 316800
804940
Less:Cost of completed production 795400
Ending balance 9540
Finished goods inventory
$
Beginning balance 34600
Add:Cost of completed production 795400
830000
Less:Cost of goods sold 801300
Ending balance 28700
Cost of goods sold=$ 801300
Allocation of over-applied overhead of $ 20540:
Ending balance % of allocation Over-applied overhead Allocated over-applied overhead
a a b a*b
Work in process inventory 9540 1.14% 20540 233
Finished goods inventory 28700 3.42% 20540 702
Cost of goods sold 801300 95.45% 20540 19605
839540 20540
Allocation of over-applied overhead will reduce Work in Process Inventory, Finished Goods Inventory, and Cost of Goods Sold balances.
Adjusted balances:
Work in process inventory=9540-233=$ 9307
Finished goods inventory=28700-702=$ 27998
Cost of goods sold=801300-19605=$ 781695